# Multi-objective Bayesian Optimisation of Spinodoid Cellular Structures for Crush Energy Absorption

### v1

@@ 5 Results
Nonetheless this work successfully demonstrates the application of the MOBO framework for crush energy absorption, highlighting its ability to balance multiple objectives and generate Pareto-optimal solutions, ultimately accelerating the selection of optimal structures. This versatile framework has the potential to be adapted for various other fields. For example, in aerospace engineering, where weight reduction, stiffness, and energy absorption are critical, it could be used to design lightweight structures tailored to specific mission profiles. Similarly, in biomedical engineering, the framework could optimise implants to maximise mechanical comparability with bone tissue while minimising material usage. In architectural design, it could assist in balancing structural integrity material efficiency, and aesthetics based on user preferences.
@@
On the optimisation side, the process could be enhanced by implementing parallel evaluation of q-batch recommendations using array jobs on an HPC cluster. This improvement would capitalise on the ability of acquisition functions to manage multiple evaluations simultaneously, thereby accelerating the optimisation process.
@@ A.1 Controlling the anisotropy of topologies
A.2 Effect of wave number on performance indicators
A.2.1 Energy Absorption
@@
A.2.1: A demonstration of the effect of varying the wave number on the EA using four test structures generated with the cone angles is shown in Fig. 2. The structures have dimensions of $40\times 40\times 40\ \text{mm}^{3}$ and were tested using different mesh resolutions, ranging from a coarse mesh resolution of 20 to a fine mesh resolution of 60. A combination of different line styles and markers is used to distinguish between the various mesh resolutions.
A.2.2 Peak Force
@@
A.2.2: Similar to A.2.1, the figure illustrates the effect of varying the wave number on the PF using the same structures.
@@
A.3 Effect of mesh resolution on stress-strain behaviour of spinodoid structures
@@
A mesh convergence study assessed its impact on the stress-strain behaviour of the four tested spinodoid cases. The mesh quality was controlled by varying the mesh resolution, which consequently altered the number of tetrahedral elements used to create the cellular structures. An increase in mesh resolution corresponds to a higher number of elements. Figure A.3.1 illustrates the stress-strain curves for various structures generated at different mesh resolutions, ranging from 20 to 60. Resolutions below 20 result in inadequately generated structures, while resolutions above 60 lead to excessively long simulation times. While the stress-strain behaviours of the structures are generally similar across this range, a closer inspection reveals that a mesh resolution of 20 exhibits the most significant difference. Further increases in mesh resolution result in minimal variations in behaviour. Consequently, a mesh resolution of 30 was selected for conducting the optimisation.
@@
A.3.1: The subplots highlight the effect of mesh resolution on the stress-strain behaviour of four spinodoid test cases generated with angles described in Fig. 2, measured at a wave number of $16\pi$. The structures were generated at various mesh resolutions, ranging from a coarse mesh resolution of 20 to a fine mesh resolution of 60.
A.4 Effect of number of samples on the Sobol’ sensitivity indices
A.4.1 Sensitivity analysis with four input variables
@@
A.4.1: Illustration showing the convergence of sensitivity indices S1 and ST as the number of samples used to calculate the Sobol’ sensitivity indices increases, with a maximum of 474 evaluations. The analysis is demonstrated for the performance indicators EA and PF. The Sobol’ sensitivity analysis was conducted assuming four input parameters ($\rho,\theta_{1},\theta_{2},\theta_{3}$).
A.4.2 Sensitivity analysis with three input variables
@@
A.4.2: Similar to A.4.1, an additional Sobol’ sensitivity study was conducted with three input parameters ($\theta_{1},\theta_{2},\theta_{3}$), with a total number of samples used in calculations reaching 1,500. This study assumes a constant relative density of $\rho=0.3$.
A.5 Benchmarks
A.5.1 Benchmark optimisation history
A.5.1: Optimisation histories for two benchmark problems that were solved using five methods: Random Sobol sampling, qEHVI, qNEHVI, weighted sum, and Chebyschev scalarisation. For the Branin-Currin problem with artificial observation noise with noise standard deviation of 15.19, and 0.63 for each objective, respectively, the covariance matrix was calculated using (a) the RBF kernel and (b) the Matern kernel. Additionally, subplots (c) and (d) illustrate the solution of the ZDT1 problem with zero additional noise, where the covariance matrix was computed using RBF and Matern kernels, respectively. The red solid line represents the analytical solution for the ZDT1 problem. The colour gradients indicate at which each specific point was obtained.
A.5.2 Benchmark cumulative hypervolume
A.5.2: The subplots present the hypervolume history over 200 iterations for two benchmark problems: Branin-Currin and ZDT1. The solid lines represent the median hypervolume, while the shaded regions above and below indicate the upper and lower bound of the hypervolumes, respectively. Subplot (a) displays the difference in hypervolumes obtained using MOBO methods compared to the NSGA-II method for the Branin-Currin problem, with the RBF kernel used for the covariance matrix. Subplot (b) shows the hypervolume difference for the Branin-Currin problem, using the Matern kernel for the covariance matrix. (c) An illustration of the hypervolume obtained from various MOBO methods for ZDT1 problem using the RBF kernel function, while (d) shows the corresponding increase in hypervolume for the ZDT1 problem using the Matern kernel. Both (c), and (d) shows the result for the first 50 iterations as beyond this the hypervolume values do not change. The reference point for the Branon-Currin, and the ZDT1 were $(18,6)$, and $(11,11)$, respectively.
A.6 Pair plot of initial dataset
A.6.1: Visualisation of the initial dataset containing 50 data points used to initialise the GP using a pair plot.
A.7 Optimisation history
A.7.1: Optimisation history for all methods implemented using the proposed optimisation framework, showing the results of the full 250 iterations and the initial data used to initialise the GP model. The left column shows the scaled outputs, whereas the right column shows the unscaled outputs, with the scaling being done according to Eq. (27). The colour bar on the right of each subplot shows the iteration at which the particular output was obtained.
A.8 HPC hardware specifications

### v2

@@ 7 Discussion
This work successfully demonstrates the application of the MOBO framework for crush energy absorption, highlighting its ability to balance multiple objectives and generate Pareto-optimal solutions, ultimately accelerating the selection of optimal structures. This versatile framework has the potential to be adapted for various other fields. For example, in aerospace engineering, where weight reduction, stiffness, and energy absorption are critical, it could be used to design lightweight structures tailored to specific mission profiles. Similarly, in biomedical engineering, the framework could optimise implants to maximise mechanical comparability with bone tissue while minimising material usage. In architectural design, it could assist in balancing structural integrity material efficiency, and aesthetics based on user preferences.
@@
On the optimisation side, the MOBO framework can be significantly enhanced by incorporating a more efficient sampling strategy that explicitly accounts for uncertainties observed in the experimental and FEM simulations. By acknowledging the inherent uncertainties in regions with large errors or inconsistent results (such as Columnar or Lamellar structures), these areas can be targetted with denser sampling. This approach would allow for more accurate predictions of key metrics—such as energy absorption and peak force—where the uncertainty is higher, while reducing the computational cost in regions where the model already performs well (e.g., Isotropic structures). This strategy aligns with the broader concept of uncertainty-aware optimisation, where sampling density is adaptively adjusted based on the observed uncertainty in the predictions. By focusing computational resources on regions with higher uncertainty, the optimisation process becomes more robust and efficient, potentially leading to more precise solutions in fewer iterations.
Furthermore, the optimisation process can be further accelerated by implementing parallel evaluation of multiple q-batch recommendations using array jobs on an HPC cluster. This would leverage the parallelism of acquisition functions, which can manage several evaluations simultaneously. By running multiple simulations or experiments in parallel, different regions of the design space can be explored at the same time, drastically reducing the time required to converge to optimal solutions. This parallelisation would not only accelerate the overall optimisation but also allows for more comprehensive exploration of the design space, especially when combined with the targeted sampling of uncertain regions. The combination of efficient sampling based on uncertainty and parallel evaluation of batch recommendations would lead to a much faster and more accurate optimisation process. Additionally, integrating these improvements within a multi-fidelity optimisation framework—which uses both experimental data and FEM simulations of varying accuracy levels—could further increase the robustness and efficiency of the process. Lower-fidelity simulations could be used in combination with high-fidelity experiments to guide the search for optimal solutions, balancing the trade-off between computational cost and model accuracy.
@@ A.1 Controlling the anisotropy of topologies
A.2 Comparison of material model with material characterisation test
To compare the FEM simulations with the material characterisation tests, three material models were utilised. The simplest was the isotropic elasticity model, which considered only linear elastic behaviour. Simulations were also conducted using an isotropic elastic-plastic model, incorporating plastic strain to simulate the post-yield behaviour of the cube under compression. Finally, the Hill anisotropic model was tested against these conventional models. This model accounts for anisotropy arising from the fabrication process while also capturing the post-yield behaviour. The same FE analysis setup was used for both the initial validation test and the one used to model the behaviour of spinodoids under compressive loading.
The comparison between the results from the FEM analyses with various material models and the experimental data revealed slight differences in the initial elastic region of the stress-strain curve. However, in the plastic deformation region, the Hill anisotropic model closely followed the experimental data, outperforming the isotropic elastic-plastic model. As the majority of energy absorption occurs within this plastic region, it is crucial to construct a FEM model that accurately replicates this behaviour.
The observed discrepancies between the Hill model and the experimental results can be attributed to multiple factors. First, the simplicity in calculating the coefficients for Hill’s anisotropic yield criterion might have contributed to the mismatch. These coefficients were derived from limited experimental data, which may not fully capture the complex anisotropic properties of the material. Additionally, the assumptions made during the formulation of the yield stress ratios-such as uniform anisotropy or idealised deformation modes-might not align perfectly with the actual material behaviour. Variability introduced during the fabrication process, such as microstructural heterogeneities, residual stresses, or variations in layer bonding during additive manufacturing, can also lead to deviations.
To reduce these discrepancies, more sophisticated calibration techniques, such as inverse modelling or optimisation-based fitting, could be employed to refine the Hill model parameters. Incorporating additional experimental tests, such as biaxial or shear tests, would also provide a more complete representation of the material’s anisotropic behaviour. Finally, accounting for fabrication-induced effects in the FEM model, such as layer-specific anisotropy or residual stress distributions, could improve agreement between the simulation and experiment.
A.2.1: Comparison of the mechanical response between the compressive testing conducted on a 3D printed cube of dimensions $20\times 20\times 20$ mm3, which was fabricated at $90^{\circ}$ orientation to the loading direction (shown by solid black line), and an equivalent FE model, which was setup with various constitutive material models (shown by various red coloured lines).
A.3 Effect of wave number on performance indicators
A.3.1 Energy Absorption
@@
A.3.1: A demonstration of the effect of varying the wave number on the EA using four test structures generated with the cone angles is shown in Fig. 2. The structures have dimensions of $40\times 40\times 40\ \text{mm}^{3}$ and were tested using different mesh resolutions, ranging from a coarse mesh resolution of 20 to a fine mesh resolution of 60. A combination of different line styles and markers is used to distinguish between the various mesh resolutions.
A.3.2 Peak Force
@@
A.3.2: Similar to A.3.1, the figure illustrates the effect of varying the wave number on the PF using the same structures.
A.4 Effect of mesh resolution on the number of elements
A.4.1 illustrates the effect of mesh resolution on the number of tetrahedral elements used to discretise a structure with the parameters $\Theta=(0.3,15\pi,15,15,0)$. While the number of elements presented corresponds specifically to this structure, it is important to note that the element count will vary depending on the resolution and the geometry of the structure.
Resolution Number of elements Average element size (mm)
20 18728 0.4398
30 84284 0.2731
40 253485 0.2002
50 526626 0.1583
60 932918 0.1316
A.4.1: This example shows how the resolution changes the number of linear tetrahedral elements of type C3D4 for a structure with parameters $\Theta=(0.3,15\pi,15,15,0)$, with the average element size shown through the equivalent length found by dividing the element volume by its area.
@@ A.4 Effect of mesh resolution on the number of elements
A.5 Effect of mesh resolution on stress-strain behaviour of spinodoid structures
@@
A mesh convergence study assessed the impact of reesolution on the stress-strain behaviour of the four tested spinodoid cases. The mesh quality was controlled by varying the mesh resolution, which consequently altered the number of tetrahedral elements used to create the cellular structures. An increase in mesh resolution corresponds to a higher number of elements. A.5.1 illustrates the stress-strain curves for various structures generated at different mesh resolutions, ranging from 20 to 60. Resolutions below 20 result in inadequately generated structures, while resolutions above 60 lead to excessively long simulation times. While the stress-strain behaviours of the structures are generally similar across this range, a resolution of 20 exhibited the most pronounced difference compared to higher resolutions. Beyond a resolution of 30, the variations in behaviour were negligible. Therefore, a resolution of 30 was chosen for the optimisation study as it provides a balance between computational efficiency and accuracy.
@@
A.5.1: The subplots highlight the effect of mesh resolution on the stress-strain behaviour of four spinodoid test cases generated with angles described in Fig. 2, measured at a wave number of $16\pi$. The structures were generated at various mesh resolutions, ranging from a coarse mesh resolution of 20 to a fine mesh resolution of 60.
A.6 Effect of time scaling and mass scaling on stress-strain behaviour
To reduce the solution time for ABAQUS dynamic Explicit simulations, there exists two common methods: Time scaling, and mass scaling. ABAQUS Explicit requires stable time increment $\Delta t=L^{e}/C_{d}$ to obtain a stable solution, where $L^{e}$ is the characteristic element length defined by the smallest element used to mesh the geometry, and $C_{d}=\sqrt{E/\rho}$ is the wave propagation speed, with $E$ being the Young’s modulus and density of the material, respectively. For quasi-static speed at which the spinodoids were tested, the stable time dictates the number of increments required to solve the problem. Time scaling allows a reduction of the number of increments required as number of increments = $t/\Delta t$, where $t$ is the total time period. In this case, the computation time reduces as much as the time was reduced by artificially increasing the velocity. In contrast, the stable time increment can also be increased, which is called mass scaling, whereby the density is artificially increased. This decreases the computation time by the square root of the density.
A.6.1 Effect of time scaling
The mechanical response of spinodoid structures was evaluated at three different total time periods: 0.001, 0.002, and 0.005 seconds. The analysis was conducted using the geometries described in Fig. 2 and compared with the corresponding experimental results, as shown in A.6.1. For the ‘Columnar’ structure, increasing the total time period led to a rapid reduction in initial stiffness, and a decrease in random fluctuations in stress values. Additionally, the time required for stress values to increase after contact between the structure and the loader measured by the stationary anvil was reduced, resulting in a better match of the small-strain response with experimental data, particularly at higher total time periods.
In the case of the ‘Isotropic’ structure, the stress-strain behaviour showed minimal variation with changes in the total time period. However, the initial stiffness became more consistent with experimental results as the total time period increased. Similar trends were observed for the ‘Cubic’ and ‘Lamellar’ structures. For the ‘Cubic’ structure, stress values at large strains began to deviate from experimental results at longer time periods. In contrast, for the ‘Lamellar’ structure, the stress-strain response at high strain values aligned more closely with experimental data as the total time period increased. Based on these observations, it was essential to balance simulation accuracy and computational efficiency. At shorter total time periods, the higher loading velocity increases kinetic energy, making inertia effects more prominent. This introduces unintended strain dependency, leading to a non-linear small-strain response. However, shorter time periods also significantly reduce computation time. To achieve an optimal balance, a total time period of 0.002 seconds was selected.
A.6.1: Plot showing the effect of time scaling for each test geometry described in Fig. 2 on the stress-strain response of each structure obtained from FEM simulations. The behaviour of each structure was tested under three different total time periods shown by the green, red, and blue lines indicating time periods of 0.001, 0.002, and 0.005 seconds, respectively, while the different black lines show the results from experiment.
A.6.2 Effect of mass scaling
A structure generated using the parameters $\Theta=(0.6,15\pi,15,15,0)$ was used to test the effects of mass scaling. Three different fixed mass scaling (FMS) values were tested, showing minimal differences in the force-displacement curves, indicating no significant loss of accuracy as illustrated in A.6.2. Increasing the FMS value led to a reduction in total simulation time, with an increase of 0.5 resulting in a 14% reduction in computation time. However, it also slightly reduced the energy absorption calculated from the force-displacement curve. Further increases in the FMS value continued to reduce simulation time, from 981 seconds at an FMS value of 1 to 636 seconds at a value of 2. Despite these performance gains, the energy absorption values showed only a partial reduction. However, artificially increasing the density (a consequence of mass scaling) reduced the wave speed, delaying the reaction force’s propagation through the structure. This delay caused an initial lag in the force-displacement curve before stress values began to rise. Ultimately, an FMS value of 2 was selected, as it provided a balance between reduced simulation time and acceptable accuracy. Further increases in FMS resulted in diminished the accuracy of the mechanical response, making higher values unsuitable.
A.6.2: Plots showing the effect of applying mass scaling to geometry with parameters $\Theta=(0.6,15\pi,15,15,0)$. The plot on the left shows the effect of three different levels of mass scaling on the mechanical response using force-displacement graph, while the plot on the right shows its effect on the simulation time and the calculated energy absorption value.
A.7 Effect of number of samples on the Sobol’ sensitivity indices
A.7.1 Sensitivity analysis with four input variables
@@
A.7.1: Illustration showing the convergence of sensitivity indices S1 and ST as the number of samples used to calculate the Sobol’ sensitivity indices increases, with a maximum of 474 evaluations. The analysis is demonstrated for the performance indicators EA and PF. The Sobol’ sensitivity analysis was conducted assuming four input parameters ($\rho,\theta_{1},\theta_{2},\theta_{3}$).
A.7.2 Sensitivity analysis with three input variables
@@
A.7.2: Similar to A.7.1, an additional Sobol’ sensitivity study was conducted with three input parameters ($\theta_{1},\theta_{2},\theta_{3}$), with a total number of samples used in calculations reaching 1,500. This study assumes a constant relative density of $\rho=0.3$.
A.8 Benchmarks
A.8.1 Benchmark optimisation history
A.8.1: Optimisation histories for two benchmark problems that were solved using five methods: Random Sobol sampling, qEHVI, qNEHVI, weighted sum, and Chebyschev scalarisation. For the Branin-Currin problem with artificial observation noise with noise standard deviation of 15.19, and 0.63 for each objective, respectively, the covariance matrix was calculated using (a) the RBF kernel and (b) the Matern kernel. Additionally, subplots (c) and (d) illustrate the solution of the ZDT1 problem with zero additional noise, where the covariance matrix was computed using RBF and Matern kernels, respectively. The red solid line represents the analytical solution for the ZDT1 problem. The colour gradients indicate at which each specific point was obtained.
A.8.2 Benchmark cumulative hypervolume
A.8.2: The subplots present the hypervolume history over 200 iterations for two benchmark problems: Branin-Currin and ZDT1. The solid lines represent the median hypervolume, while the shaded regions above and below indicate the upper and lower bound of the hypervolumes, respectively. Subplot (a) displays the difference in hypervolumes obtained using MOBO methods compared to the NSGA-II method for the Branin-Currin problem, with the RBF kernel used for the covariance matrix. Subplot (b) shows the hypervolume difference for the Branin-Currin problem, using the Matern kernel for the covariance matrix. (c) An illustration of the hypervolume obtained from various MOBO methods for ZDT1 problem using the RBF kernel function, while (d) shows the corresponding increase in hypervolume for the ZDT1 problem using the Matern kernel. Both (c), and (d) shows the result for the first 50 iterations as beyond this the hypervolume values do not change. The reference point for the Branon-Currin, and the ZDT1 were $(18,6)$, and $(11,11)$, respectively.
A.9 Pair plot of initial dataset
A.9.1: Visualisation of the initial dataset containing 50 data points used to initialise the GP using a pair plot.
A.10 Optimisation history
A.10.1: Optimisation history for all methods implemented using the proposed optimisation framework, showing the results of the full 250 iterations and the initial data used to initialise the GP model. The left column shows the scaled outputs, whereas the right column shows the unscaled outputs, with the scaling being done according to Eq. (27). The colour bar on the right of each subplot shows the iteration at which the particular output was obtained.
A.11 HPC hardware specifications
@@
A.12 Optimisation history with three variables
A.12.1: Optimisation history of optimisation task conducted with three variables, $\theta$’s, with fixed relative density at $\rho=0.3$, showing the results of full 250 iterations with 50 initial data points, conducted using the qNEHVI method. The figure on the left shows the optimisation history with normalised objectives scaled according to Eq. (27), while the figure on the right shows the history with the normalised objectives. While the colour bar shows the iteration at which the output was obtained.
A.13 Evolution of recommended design parameters
A.13.1: Left column shows the evolution of recommended design parameters for three-variable optimisation with a fixed relative density at $\rho=0.3$, whereas the right column shows the evolution for four-variable optimisation with $\rho$ varying between 0.3 and 0.6. The first 50 points shown in each row shows the initial data, while the rest show the design parameters required to be evaluated throughout the optimisation.